OpenStudy (anonymous):

5. The height of a cylinder is three times the diameter of the base. The surface area of the cylinder is 126pi ft2. What is the radius of the base?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

The height of a cylinder is three times the diameter of the base. h = 6 times the radius h = 6r Surface area = 2PI r h + 2 PI r^2 S = 2 π r^2 + 2 π r h 6.28319r^2+37.69911r-395.84034 This equation is of form ar^2+br+c a = 6.28319 b = 37.69911 c = -395.84034 r=[-b+/-sqrt(b^2-4ac)]/2a] r=[-37.69911 +/-sqrt(37.69911^2-4(6.28319)(-395.84034… discriminant is b^2-4ac =11369.775867 r=[-37.69911 +√(106.62915)] / (2)(6.28319) r=[-37.69911 -√(106.62915)] / (2)(6.28319) r=[-37.69911+106.62915] / 12.56637 r=[-37.69911-106.62915] / 12.56637 The roots are 5.4853 and -11.4853 Radius = 5.4853 Radius = -11.4853 Radius cannot be negative, so radius = 5.4853

OpenStudy (anonymous):